A cannonball is dropped from the top of a building. If the point of release is 105 ft above the ground, what is the velocity of the cannonball just before it strikes the ground?
1
Expert's answer
2019-11-13T09:28:29-0500

The law of conservation of energy gives


"mgh=\frac{mv^2}{2}"

So, the velocity of the cannonball just before it strikes the ground


"v=\sqrt{2gh}"


"=\sqrt{2\times 32\:\rm ft/s^2\times 105\: ft}=82\:\rm ft/s=25\:\rm m/s"